The perpendicular distance of point (2,-3,-4)from z axis is
 

Grade:11

1 Answers

Arun
5 years ago
Dear Yuvraj
 
The perpendicular distance of point (2,-3,-4)from z axis  = sqrt (2^2 + -3^2) = sqrt (4+9) = sqrt (13)
 
 
Regards
Arun (askIITias forum expert)